OpenStudy (anonymous):

Discs of radius 4cm are cut from a rectangular sheet of length 84cm and width 24cm. How many complete discs can be cut out?

OpenStudy (anonymous):

So a 4*4 cm square. Since the width is 24cm then we can fit 6 disks and the height is 84cm so we can fit 21 6*21=126 disks

OpenStudy (anonymous):

Area of the rectangular sheet = 2016 sq. cm Area of one disk = 50.24 sq. cm. So now you just have to divide.

OpenStudy (anonymous):

40??

OpenStudy (anonymous):

but my text gives the answer 30

OpenStudy (anonymous):

Yes, its 40.

OpenStudy (anonymous):

Textbooks can be occasionally wrong.....

OpenStudy (anonymous):

And Zed, you must have misunderstood, a disk is a circle, not a square.

OpenStudy (anonymous):

Sorry I thought the diameter was 4. So we can have 3 disk in the width (24/8) and 10 disk in the height (84/8). So the total number is 3*10=30

OpenStudy (anonymous):

Because the distance from the center of the circle is 4cm. It must fit in a 8*8cm square.

OpenStudy (anonymous):

@pratu043, the area of 40 disks can be cut out of the sheet, however they will not be the shape of a disk.

OpenStudy (anonymous):

why did you divide the diameter by the length and height of the rectangle????

OpenStudy (anonymous):

and were did the 10 come from???

OpenStudy (anonymous):

okay the disks will fit in a 8x8 square. do you understand this?

OpenStudy (anonymous):

where did you get 8x8 from?

OpenStudy (anonymous):

okay because the radius of the circle is 4, its diameter must be 8. and the diameter cannot change so it is 8 for the whole circle. Now because the circle is not longer than 8 cm it must fit into a 8x8 square. do you follow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so now what i'm trying to find is how many of these squares fit into the rectangular sheet.

OpenStudy (anonymous):

for the width i can fit 3. because 24/8=3 for the height i can fit 10. because 84/8=10.5 but 10.5 won't make a disk so we must we 10. So the the amount of disks=number for width*number for height =3*10 = 30

OpenStudy (anonymous):

impressive